OpenStudy (anonymous):

evaluate log equation help! 9^(logbase3(20))

OpenStudy (anonymous):

I don't know how to solve it in this format, could someone please tell me how these equations work?

OpenStudy (zarkon):

\[\Large9^{\log_{3}(20)}=(3^2)^{\log_{3}(20)}=3^{2\log_{3}(20)}=\left(3^{\log_{3}(20)}\right)^2\]

OpenStudy (anonymous):

what happens now that the 2 have been factored out?

OpenStudy (anonymous):

@Zarkon